calculate PH of a solution whose 100 mL contains 0.2 g NaOH dissolved in it

Grade:

2 Answers

Himanshu Rathour
5 years ago
Moles of NaOH =0.2/40
=0.005
Hence moles of OH¯=0.005
Concentration of OH¯=0.005/0.1
=0.05
pOH=-log(0.05)
1.30
pH=14-1.3
pH=12.7
Pratik parash swain
5 years ago
Molecular maa of NaOH=40
Since given 0.2 gm of NaOH is dissolved in a solution of 100 ml
Then,moles of NaOH =0.2/40=0.005=5×10-3
moles of OH -=5×10-3
concentration ofOH-=5×10-3/0.1=5×10-2
pOH=-log(5×10-2)=1.30
As we know ,pH+pOH=14
Then,
pH=14-1.30=12.7